OpenStudy (anonymous):

guys long division working for 6.4 divided by 0.04

OpenStudy (anonymous):

160

OpenStudy (anonymous):

\[6.4=64/10\] \[0.04 = 4/100\] \[6.4/0.04 = (64/10) / (4/100) = 160 \]

OpenStudy (anonymous):

None of that looks like long division...

OpenStudy (anonymous):